OpenStudy (anonymous):

Factor by grouping 6u2-13u+6

OpenStudy (hoblos):

= (3u-2)(2u-3)

OpenStudy (hoblos):

6u2-13u+6 = 6u2- 9u -4u +6 = 3u(2u-3) - 2(2u-3) = (2u-3)(3u-2)